# Measurement as Bricolage: Examining How Data Scientists Construct Target Variables for Predictive Modeling Tasks

### v2

@@ Abstract.
Data scientists often formulate predictive modeling tasks involving fuzzy, hard-to-define concepts, such as the “authenticity” of student writing or the “healthcare need” of a patient. Yet the process by which data scientists translate fuzzy concepts into a concrete, proxy target variable remains poorly understood. We interview fifteen data scientists in education (N=8) and healthcare (N=7) to understand how they construct target variables for predictive modeling tasks. Our findings suggest that data scientists construct target variables through a bricolage process, involving iterative negotiation between high-level measurement objectives and low-level practical constraints. Data scientists attempt to satisfy five major criteria for a target variable through bricolage: validity, simplicity, predictability, portability, and resource requirements. To achieve this, data scientists adaptively use problem (re)formulation strategies, such as swapping out one candidate target variable for another when the first fails to meet certain criteria (e.g., predictability), or composing multiple outcomes into a single target variable to capture a more holistic set of modeling objectives. Based on our findings, we present opportunities for future HCI, CSCW, and ML research to better support the art and science of target variable construction.
@@
copyright: ccjournal: PACMHCIjournalyear: 2025journalvolume: 9journalnumber: 7article: CSCW447publicationmonth: 11doi: 10.1145/3757628ccs: Human-centered computing Empirical studies in HCIccs: Computing methodologies Machine learningccs: Computing methodologies Supervised learning
@@ 1. Introduction
However, the traditional top-down conceptualization of measurement is in tension with our understanding of data science as a bottom-up process constrained by existing data (Passi and Barocas, 2019; Kross and Guo, 2021). Passi and Barocas (2019) observe that problem formulation in data science is an iterative process, where data scientists progressively reconcile their high-level objectives with low-level data and modeling constraints. As a result, while a traditional top-down measurement approach assumes that an analyst has the resources to collect data tailor-made to the goals of a study, data scientists are often forced to re-purpose the limited data that happens to be available. This disconnect creates a significant challenge: though many flawed data science projects can be traced back to target variable construction (e.g., (Turque, 2012; Amrein-Beardsley, 2014; Mullainathan and Obermeyer, 2021; Kawakami et al., 2022; Cheng et al., 2022)), existing tools, frameworks, and pedagogical resources reflect a top-down planning approach towards measurement, rarely accounting for the bottom-up, resource-constrained reality of data science practice.
@@
To understand data scientists’ current practices and challenges in constructing target variables—and to identify opportunities to better support target variable construction in practice—in this work, we conducted semi-structured interviews with fifteen data scientists from education (N=8) and healthcare (N=7). We asked data scientists to walk us through how they navigate target variable construction in prior modeling projects. Participants also engaged in a hypothetical target variable construction task, allowing us to observe their reasoning in a new modeling context. Our goal in these interviews was to understand how data scientists intertwine two conceptually distinct activities, measurement of latent constructs and prediction of target variables, during problem formulation (Mussgnug, 2022). While prior empirical studies of data science practice have often centered prediction (e.g., (Alspaugh et al., 2018; Wongsuphasawat et al., 2019; Kandel et al., 2012)), we broaden our scope of inquiry to investigate the interplay between measurement validity and predictive utility in problem formulation.
@@
We find that data scientists construct target variables by leveraging creative and pragmatic approaches to make do with the limited data they have at hand. Data scientists craft their formulation through a process of iterative dialogue with their data, tweaking and refining their target variable until they conclude that it provides a sufficient operationalization for the unobserved construct of interest. To understand data scientists’ target variable construction practices, we draw upon the concept of bricolage, originally introduced by Claude Lévi-Strauss to contrast the rigid top-down methodology of Western science with the more tangible “science of the concrete” practiced in non-Western societies (Levi-Strauss, 1966). Rather than implementing abstract concepts through top-down design, the bricoleur solves problems by arranging and re-arranging well-known materials in a bottom-up process. Like bricoleurs, data scientists in our study often acknowledged imperfection in their improvisational approach to measurement, yet they also identified innovative solutions that would be difficult to anticipate through a process of top-down design.
@@ 2. Background and Related Work
In this section, we begin by introducing Claude Lévi-Strauss’ theory of Bricolage. We then describe existing Science and Technology Studies (STS), HCI, and CSCW research that has investigated questions related to categorization, measurement, and target variable construction in data science practice. We conclude by providing an overview of key concepts from measurement theory in the quantitative social sciences which we build upon in our interviews, findings, and discussion.
@@ 2.1. Bricolage and the Science of the Concrete
Bricolage, defined by Claude Lévi-Strauss as “making the most of available resources”(Levi-Strauss, 1966), describes how actors (i.e., bricoleurs) construct artifacts in creative and adaptive ways. Bricolage is central to what Lévi-Strauss calls the science of the concrete, which contrasts with modern scientific thought by focusing on the materials directly at hand rather than abstract principles. Bricolage contains three elements: repertoire, dialogue, outcome (Duymedjian and Rüling, 2010). Repertoire describes the resources—physical artifacts, ideas, or skills—that the bricoleur brings to bear on a problem. Whereas the engineer or scientist acquires resources tailor-made to specification, the bricoleur assembles “odds and ends” that might be useful in the future. Then, when faced with a new problem, the bricoleur engages in dialogue with these materials—a process of arranging and re-arranging materials in her repertoire to find a workable solution. While engaging in dialogue, the bricoleur adopts old objects for new purposes and combines concepts with materials in novel ways. Finally, outcome describes the artifact resulting from the bricolage process. While bricolage entails a tolerance for imperfection, bricoleurs sometimes identify “brilliant unforeseen results” (Levi-Strauss, 1966) as part of their process.
@@
In this work, we draw a novel connection between data science work and bricolage. Much like an entrepreneur, programmer, or painter, data scientists construct target variables by combining existing resources in creative, adaptive ways. Instead of instantiating abstract plans through top-down design—akin to the structured programming approach taught in classrooms—data scientists engage in more direct dialogue with their data and modeling tools, making do with the materials at hand. And analogous to past research investigating how to support such exploratory and improvisational coding practices (Kery and Myers, 2017), this has implications for the design of practical tools to support target variable construction in practice. We highlight connections between target variable construction and bricolage throughout the findings ($\S$ 4) and describe implications for tool design in the discussion ($\S$ 5).
@@ 2.2. Tracing the Social Construction of Target Variables
Target variable construction also relates to a rich body of Science and Technology Studies (STS) and philosophy of science literature that investigates the social processes that give rise to categories in data (Latour and Woolgar, 2013; Bowker, 2000; Pine and Liboiron, 2015; Harvey et al., 2024; Porter, 1996; Crosby, 1997; Hacking, 1999; Mol, 2002; Suchman, 1993). Latour and Woolgar (2013) describe the social construction of scientific facts within laboratories, highlighting that facts are not merely discovered but actively produced through a complex interplay of social and technical factors. Likewise, Bowker (2000) examines how the International Classification of Diseases (ICD), shape and are shaped by the social and political contexts in which they are embedded. Building on Bowker’s work, recent studies have also explored the construction of racial and ethnic categories in AI training data (e.g., (Abdu et al., 2023; Mickel, 2024)) and charted measurement assumptions underpinning motion capture technologies (Harvey et al., 2024).
@@ 2.3. Empirical Studies of Data Science Practice
Our work also builds upon empirical studies investigating how data scientists conduct their day-to-day work. Many studies have examined activities in the technical “inner loop” (Kross and Guo, 2021) of data science practice, such as exploratory data analysis (e.g., (Alspaugh et al., 2018; Wongsuphasawat et al., 2019)), feature engineering (e.g., (Muller et al., 2019)), and visualization (e.g., (Kandel et al., 2012)). Other studies have investigated broader “outer loop” activities performed by data scientists, including collaborating with other stakeholders (Kross and Guo, 2021; Passi and Jackson, 2018; Nahar et al., 2022; Mao et al., 2019; Kim et al., 2016; Zhang et al., 2020; Hou and Wang, 2017) and addressing ethical concerns (Holstein et al., 2019; Deng et al., 2022). CSCW research has described inner and outer loop activities as iterative and interconnected (Kross and Guo, 2021; Passi and Barocas, 2019). In particular, based on interviews with ten professionals spanning industry and academia, Kross and Guo (2021) observed that data scientists iteratively “framed the problem” (i.e., an outer loop activity) and returned to their data to gain an understanding of its affordances (i.e., inner loop activities).
However, we currently have an incomplete understanding of how data scientists construct target variables while formulating prediction tasks. To date, Passi and Barocas (2019) provide the most detailed account of how problem formulation unfolds in real-world modeling contexts. Over the course of an eighteen-month ethnographic study of a corporate data science team, Passi and Barocas (2019) observed that data scientists repeatedly (re)formulated their prediction task as they discovered new information about data availability constraints and the predictability of alternative outcome variables. Based on their findings, Passi and Barocas (2019) describe problem formulation as a negotiated translation between high-level objectives and a concrete prediction task. Our work builds upon this finding to identify the more general evaluation desiderata and (re)formulation strategies driving the problem formulation process. We develop a process model through analysis of our findings and use this model to identify targeted opportunities to better support data scientists’ problem formulation practices going forward. For example, when viewed through the lens of our findings, the team studied by Passi and Barocas (2019) used one of five (re)formulation strategies (i.e., swapping) to change outcome variables when the first proved infeasible due to data availability constraints. We present design opportunities to help data scientists more effectively apply swapping ($\S$ 4.2.3) and other (re)formulation strategies.
@@ 2.4. Toolkits Supporting Target Variable Construction
To mitigate these issues, the HCI, ML, and FAccT research communities have developed a range of toolkits to intervene in data scientists’ inner and outer loop workflows. For instance, researchers have developed systems facilitating data collection and configuration (e.g., (Bhattacharya et al., 2024; Guo et al., 2011; Kandel et al., 2011; Rattenbury et al., 2017)), exploratory data analysis (e.g., (Wongsuphasawat et al., 2015; Kery et al., 2018)), model training (e.g., (Pang et al., 2020; Dingen et al., 2018)), caching intermediate states during data science experiments (e.g., (Schubotz et al., 2022)), and performance assessment (e.g., (Cabrera et al., 2023a; Yan et al., 2021)). Many tools have also been developed to help data scientists explain their models (e.g., (Ribeiro et al., 2018, 2016; Lundberg, 2017)) or assess fairness-related concerns (e.g., (Tramer et al., 2017; Bellamy et al., 2019; Weerts et al., 2023; Saleiro et al., 2018; Liu et al., 2024b; Cabrera et al., 2019; Ahn and Lin, 2019; Wexler et al., 2019; Dingen et al., 2018)). In the FAccT community, researchers have also developed responsible AI checklists and guidelines to help data scientists assess and document components of their AI system (e.g., dataset information (Gebru et al., 2021; Pushkarna et al., 2022), intended use cases (Mitchell et al., 2019)). To our knowledge, Gala et al. (2024) propose the only system specifically designed to support the target variable construction process. Instead, existing toolkits are often designed to support data scientists’ practices only after a target variable has been specified. This presents a pressing gap in data science support at a critical point in their workflows.
However, prior human–computer interaction research suggests that such toolkits may prove ineffective if they are not designed around the existing practices of data scientists. For instance, past efforts to improve programming tools without a human-centered assessment of programmers’ needs yielded ineffective languages and systems (Myers et al., 2016; Hazzan and Tomayko, 2004; Seffah et al., 2005; Ko et al., 2011). Similarly, the utility of ML fairness toolkits was found to be limited in practice as they were designed for isolated use by technical roles, while actual fairness work often relied on deep cross-functional collaborations (Deng et al., 2022). Our work aims to help the research community avoid similar challenges while developing tools to support data scientists’ problem formulation processes. In particular, we identify opportunities for the HCI, CSCW, and ML research communities to develop tools that help data scientists more carefully weigh trade-offs across multiple evaluation criteria ($\S$ 5.3.1), select ($\S$ 5.3.2) and apply ($\S$ 5.3.3) strategies, and evaluate the validity of their problem formulation more effectively ($\S$ 5.3.4).
@@ 4.2. Strategies for (Re)formulating Prediction Tasks
Given their limited and imperfect data resources, how do data scientists engage in bricolage to craft a formulation that satisfies their criteria? Much like a painter refining a painting through successive brushstrokes, or a programmer compiling and re-compiling a system after adjusting lines of code (Turkle, 2011), data scientists iteratively construct target variables through (re)formulation strategies. These strategies form the backbone of the dialogue between the data scientist and their formulation, enabling them to respond adaptively to constraints encountered while working with existing data. At times, data scientists apply these strategies instinctively, without a second thought. At other times, encountered eureka moments after discovering a strategy that leveraged unexpected data sources or combined existing ones in novel ways. In the following sections, we detail five key strategies—piggybacking, composing, swapping, bridging, and refining—that data scientists used to navigate the constraints of their limited repertoire.

### v3

@@ Abstract.
Data scientists often formulate predictive modeling tasks involving fuzzy, hard-to-define concepts, such as the “authenticity” of student writing or the “healthcare need” of a patient. Yet the process by which data scientists translate fuzzy concepts into a concrete, proxy target variable remains poorly understood. We interview fifteen data scientists in education (N=8) and healthcare (N=7) to understand how they construct target variables for predictive modeling tasks. Our findings suggest that data scientists construct target variables through a bricolage process, in which they use creative and pragmatic approaches to make do with the limited data at hand. Data scientists attempt to satisfy five major criteria for a target variable through bricolage: validity, simplicity, predictability, portability, and resource requirements. To achieve this, data scientists adaptively apply problem (re)formulation strategies, such as swapping out one candidate target variable for another when the first fails to meet certain criteria (e.g., predictability), or composing multiple outcomes into a single target variable to capture a more holistic set of modeling objectives. Based on our findings, we present opportunities for future HCI, CSCW, and ML research to better support the art and science of target variable construction.
@@
copyright: ccjournal: PACMHCIjournalyear: 2025journalvolume: 9journalnumber: 7article: CSCW447publicationmonth: 11doi: 10.1145/3757628isbn: 978-1-4503-XXXX-X/2018/06ccs: Human-centered computing Empirical studies in HCIccs: Computing methodologies Machine learningccs: Computing methodologies Supervised learning
@@ 1. Introduction
However, the traditional top-down conceptualization of measurement is in tension with our understanding of data science as a bottom-up process constrained by existing data (Passi and Barocas, 2019; Kross and Guo, 2021). Passi and Barocas (2019) observe that problem formulation in data science is an iterative process, where data scientists progressively reconcile their high-level objectives with low-level data and modeling constraints. As a result, while a traditional top-down measurement approach assumes that an analyst has the resources to collect data tailor-made to the goals of a study, data scientists are often forced to re-purpose the limited data that happens to be available. This disconnect creates a significant challenge: though many flaws in data science projects can be traced back to target variable construction (e.g., (Turque, 2012; Amrein-Beardsley, 2014; Mullainathan and Obermeyer, 2021; Kawakami et al., 2022; Cheng et al., 2022)), existing tools, frameworks, and pedagogical resources reflect a top-down planning approach towards measurement, rarely accounting for the bottom-up, resource-constrained reality of data science practice.
@@
To understand data scientists’ current practices and challenges in constructing target variables—and to identify opportunities to better support target variable construction in practice—in this work, we conducted semi-structured interviews with fifteen data scientists from education (N=8) and healthcare (N=7). We asked data scientists to walk us through how they navigated target variable construction in prior modeling projects. Participants also engaged in a hypothetical target variable construction task, allowing us to observe their reasoning in a new modeling context. Our goal in these interviews was to understand how data scientists intertwine two conceptually distinct activities, measurement of latent constructs and prediction of target variables, during problem formulation (Mussgnug, 2022). While prior empirical studies of data science practice have often centered prediction (e.g., (Alspaugh et al., 2018; Wongsuphasawat et al., 2019; Kandel et al., 2012)), we broaden our scope of inquiry to investigate the interplay between measurement validity and predictive utility in problem formulation.
@@
We find that data scientists construct target variables by leveraging creative and pragmatic approaches to make do with the limited data they have at hand. Data scientists craft their formulation through a process of iterative dialogue with their data, tweaking and refining their target variable until they conclude that it provides a sufficient operationalization for the unobserved construct of interest. To understand data scientists’ target variable construction practices, we draw upon the concept of bricolage, originally introduced by Claude Lévi-Strauss to contrast the rigid top-down methodology of Western science with the more tangible “science of the concrete” practiced in non-Western societies (Lévi-Strauss, 1966). Rather than implementing abstract concepts through top-down design, the bricoleur solves problems by arranging and re-arranging well-known materials in a bottom-up process. Like bricoleurs, data scientists in our study often acknowledged imperfection in their improvisational approach to measurement. Yet they also identified innovative solutions that would be difficult to anticipate through a process of top-down design.
@@ 2. Background and Related Work
In this section, we begin by introducing Claude Lévi-Strauss’ theory of bricolage. We then describe existing Science and Technology Studies (STS), HCI, and CSCW research that has investigated questions related to categorization, measurement, and target variable construction in data science practice. We conclude by providing an overview of key concepts from measurement theory in the quantitative social sciences which we build upon in this work.
@@ 2.1. Bricolage and the Science of the Concrete
Bricolage, defined by Claude Lévi-Strauss as “making the most of available resources”(Lévi-Strauss, 1966), describes how actors (i.e., bricoleurs) construct artifacts in creative and adaptive ways. Bricolage is central to what Lévi-Strauss calls the science of the concrete, which contrasts with modern scientific thought by focusing on the materials directly at hand rather than abstract principles. Bricolage contains three elements: repertoire, dialogue, outcome (Duymedjian and Rüling, 2010). Repertoire describes the resources—physical artifacts, ideas, or skills—that the bricoleur brings to bear on a problem. Whereas the engineer or scientist acquires resources tailor-made to specification, the bricoleur assembles “odds and ends” that might be useful in the future. Then, when faced with a new problem, the bricoleur engages in dialogue with these materials—a process of arranging and re-arranging materials in her repertoire to find a workable solution. While engaging in dialogue, the bricoleur adopts old objects for new purposes and combines concepts with materials in novel ways. Finally, outcome describes the artifact resulting from the bricolage process. While bricolage entails a tolerance for imperfection, bricoleurs sometimes identify “brilliant unforeseen results” (Lévi-Strauss, 1966) as part of their process.
@@
In this work, we draw a novel connection between data science work and bricolage. Much like an entrepreneur, programmer, or painter, data scientists construct target variables by combining existing resources in creative, adaptive ways. Instead of instantiating abstract plans through top-down design—akin to the structured programming approach taught in classrooms—data scientists engage in more direct dialogue with their data and modeling tools, making do with the materials at hand. And analogous to past research investigating how to support such exploratory and improvisational coding practices (Kery and Myers, 2017), this has implications for the design of tools to support target variable construction in practice. We highlight connections between target variable construction and bricolage throughout the findings ($\S$ 4) and describe implications for tool design in the discussion ($\S$ 5).
@@ 2.2. Tracing the Social Construction of Target Variables
Target variable construction also relates to a rich body of Science and Technology Studies and philosophy of science literature that investigates the social processes that give rise to categories in data (Latour and Woolgar, 2013; Bowker, 2000; Pine and Liboiron, 2015; Harvey et al., 2024; Porter, 1996; Crosby, 1997; Hacking, 1999; Mol, 2002; Suchman, 1993). Latour and Woolgar (2013) describe the social construction of scientific facts within laboratories, highlighting that facts are not merely discovered but actively produced through a complex interplay of social and technical factors. Likewise, Bowker (2000) examines how the International Classification of Diseases (ICD), shape and are shaped by the social and political contexts in which they are embedded. Building on Bowker’s work, recent studies have also explored the construction of racial and ethnic categories in AI training data (e.g., (Abdu et al., 2023; Mickel, 2024)) and charted measurement assumptions underpinning motion capture technologies (Harvey et al., 2024).
@@ 2.3. Empirical Studies of Data Science Practice
Our work also builds upon empirical studies investigating how data scientists conduct their day-to-day work. Many studies have examined activities in the technical “inner loop” (Kross and Guo, 2021) of data science, such as exploratory data analysis (Alspaugh et al., 2018; Wongsuphasawat et al., 2019), feature engineering, (Muller et al., 2019), and visualization, (Kandel et al., 2012). Other studies have investigated “outer loop” activities performed by data scientists, including collaborating with other stakeholders (Kross and Guo, 2021; Passi and Jackson, 2018; Nahar et al., 2022; Mao et al., 2019; Kim et al., 2016; Zhang et al., 2020; Hou and Wang, 2017) and addressing ethical concerns (Holstein et al., 2019; Deng et al., 2022). Prior work has characterized inner and outer loop activities as iterative and interconnected (Kross and Guo, 2021; Passi and Barocas, 2019). In particular, based on interviews with ten professionals spanning industry and academia, Kross and Guo (2021) observed that data scientists iteratively “framed the problem” (i.e., an outer loop activity) then returned to their data to perform analyses (i.e., inner loop activities).
However, we currently have an incomplete understanding of how data scientists construct target variables while formulating prediction tasks. To date, Passi and Barocas (2019) provide the most detailed account of how problem formulation unfolds in real-world modeling contexts. Over the course of an eighteen-month ethnographic study of a corporate data science team, Passi and Barocas (2019) observed that data scientists repeatedly (re)formulated their prediction task as they discovered new information about data availability constraints and the predictability of alternative outcome variables. Based on their findings, Passi and Barocas (2019) describe problem formulation as a negotiated translation between high-level objectives and a concrete prediction task. Our work builds upon this finding to identify the more general evaluation desiderata and (re)formulation strategies driving the problem formulation process. We develop a process model through analysis of our findings and use this model to identify targeted opportunities to better support data scientists’ problem formulation practices going forward. For example, when viewed through the lens of our findings, the team studied by Passi and Barocas (2019) used one of five (re)formulation strategies (i.e., swapping) to change outcome variables when the first proved infeasible due to data availability constraints. We present design opportunities to help data scientists more effectively apply swapping and other (re)formulation strategies in Section 5.
@@ 2.4. Toolkits Supporting Target Variable Construction
To mitigate these issues, the HCI, ML, and FAccT research communities have developed a range of toolkits to intervene in data scientists’ inner and outer loop workflows. For instance, researchers have developed systems facilitating data collection and configuration (e.g., (Bhattacharya et al., 2024; Guo et al., 2011; Kandel et al., 2011; Rattenbury et al., 2017)), exploratory data analysis (e.g., (Wongsuphasawat et al., 2015; Kery et al., 2018)), model training (e.g., (Pang et al., 2020; Dingen et al., 2018)), caching intermediate states during data science experiments (e.g., (Schubotz et al., 2022)), and performance assessment (e.g., (Cabrera et al., 2023a; Yan et al., 2021)). Many tools have also been developed to help data scientists explain their models (e.g., (Ribeiro et al., 2018, 2016; Lundberg, 2017)) or assess fairness-related concerns (e.g., (Tramer et al., 2017; Bellamy et al., 2019; Weerts et al., 2023; Saleiro et al., 2018; Liu et al., 2024b; Cabrera et al., 2019; Ahn and Lin, 2019; Wexler et al., 2019; Dingen et al., 2018)). In the FAccT community, researchers have also developed responsible AI checklists and guidelines to help data scientists assess and document components of their AI system (e.g., dataset information (Gebru et al., 2021; Pushkarna et al., 2022), intended use cases (Mitchell et al., 2019)). To our knowledge, Gala et al. (2024) and Sivaraman et al. (2025) propose the only systems specifically designed to support target variable construction. Instead, existing toolkits are often designed to support data scientists’ practices only after a target variable has been specified. This presents a pressing gap in data science support at a critical point in their workflows.
However, prior human–computer interaction research suggests that new toolkits may prove ineffective if they are not designed around the existing practices of data scientists. For instance, past efforts to improve programming tools without a human-centered assessment of programmers’ needs yielded ineffective languages and systems (Myers et al., 2016; Hazzan and Tomayko, 2004; Seffah et al., 2005; Ko et al., 2011). Similarly, the utility of ML fairness toolkits was found to be limited in practice as they were designed for isolated use by technical roles, while actual fairness work often relied on deep cross-functional collaborations (Deng et al., 2022). Our work aims to help the research community avoid similar challenges while developing tools to support data scientists’ problem formulation processes. In particular, we identify opportunities for the HCI, CSCW, and ML research communities to develop tools that help data scientists more carefully weigh trade-offs across multiple evaluation criteria ($\S$ 5.3.1), select ($\S$ 5.3.2) and apply ($\S$ 5.3.3) strategies, and evaluate the validity of their problem formulation more effectively ($\S$ 5.3.4).
@@ 4.2. Strategies for (Re)formulating Prediction Tasks
Given their limited and imperfect data resources, how do data scientists engage in bricolage to craft a formulation that satisfies their criteria? Much like a painter refining a painting through successive brushstrokes, or a programmer compiling and re-compiling a system after adjusting lines of code (Turkle, 2011), data scientists iteratively construct target variables through (re)formulation strategies. These strategies form the backbone of the dialogue between the data scientist and their formulation, enabling them to respond adaptively to constraints encountered while working with existing data. At times, data scientists apply these strategies instinctively, without a second thought. At other times, data scientists encountered eureka moments after discovering a strategy that leveraged unexpected data sources or combined existing ones in novel ways. In the following sections, we detail five key strategies—piggybacking, composing, swapping, bridging, and refining—that data scientists used to navigate the constraints of their limited repertoire.